A recipe for chocolate cookies includes the following: 550 mL flour, 5.0 mL baking soda, 250 mL butter, 350 mL sugar (half of which is brown sugar), 5.0 mL vanilla extract, 2 eggs, 675 g of chocolate chips and 250 mL of chopped nuts. The recipe is set to make 100 cookies.


1
Expert's answer
2021-02-05T04:07:57-0500

For 100 cookies

Need:

1)55000 mL flour or 55 L

2)500 mL baking soda or 0,5 L

3)25000 mL butter or 25 L

4) 35000 mL sugar or 35 L

5)500 mL vanilla or 0,5 L

6) 200 eggs

7) 67500 g of chocolate or 67,5 kg

8) 25000 mL of chopped nuts

Sum:

1) 1--- 550

100--- X. a)X*1=550*100 b)X=55000/1 c)x=55000 (mL)

If you want I can continue.